i dont get how to set the problem up. once i get the problem set up, i can do it. please help?
Your chemistry teacher asks you to mix 20% acid solution with 50% acid solution to obtain 12 ounces of 30% acid solution. how much of each must you use?

Answer by Paul(988)

20x+50(12-x)=30(12)
20x+600-50x=360
-30x=-240
x=8

12-8=4

Hence, for 20% there should be 8 ounces added and for 50% there should be 4 ounces added.
Paul.